PapiChulo:

Show that the sum of the first 2n natural numbers is n(2n+1).

justjm:

I figured it out, let me type it out

PapiChulo:

hurry up

justjm:

The sum of natural numbers can be considered as \(\text{1+2+3+4+5...}\) We can indicate that as \(\sum_{n=0}^{∞} n+1\) It is asking for the sum of the first 2n natural numbers, so it is asking for us to prove \(\sum_{n=0}^{2n} n+1 = n(2n+1)\) By the general formula for the sum of an arithmetic series, which is \(S_{n}=\frac{n}{2}\left(2a+\left(n-1\right)d\right)\), we can now substitute the values from \(\sum_{n=0}^{2n} n+1\) \(\left(S_{2n}=\frac{2n}{2}\left(2\left(1\right)\right)+\left(2n-1\right)\left(1\right)\right)\) Simplified, that is \(S_{2n}=n\left(2+2n-1\right)=n\left(2n+1\right)\) Hence, we have proved that \(\sum_{n=0}^{2n}n+1=n\left(2n+1\right)\)

justjm:

Do you understand this now?

PapiChulo:

no your dumb

justjm:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @PapiChulo no your dumb \(\color{#0cbb34}{\text{End of Quote}}\) The fact you can't use *your and you're* correctly makes you the dumber one buddy

justus:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @PapiChulo no your dumb \(\color{#0cbb34}{\text{End of Quote}}\) He wasted his precious time on answering your question with out a flaw and gave you an explanation and this is what you have to say..shameful.